OpenStudy (anonymous):

What is the equation of a line passing through (-3, 7) and having a slope of -1/5?

OpenStudy (anonymous):

A. y = -5x + 22 B. y = 5x + 22 C. y=1/5x +32/5 D.y=-1/5x+32/5 E. y=-1/5x+22

OpenStudy (anonymous):

@AntarAzri